OpenStudy (anonymous):

Factoring a square root... sqr(x-2)-sqr(2x+6)-1 I need to turn this into a quadratic equation.

OpenStudy (anonymous):

\[\sqrt{x-2}-\sqrt{2x+6}-1\]

OpenStudy (anonymous):

Square everything and solve normally.

OpenStudy (anonymous):

ACK! That was so simple... I don't know why I didn't remember that!!!!

OpenStudy (anonymous):

Now would the negative signs turn to positives?

OpenStudy (anonymous):

I mean the central negative sign and the sign in front of 1

OpenStudy (anonymous):

Nope. All of the quantities are squared. Like, each of the roots and it's -(1^2) for the constant.

OpenStudy (anonymous):

Okay great, that helped out so much!

OpenStudy (anonymous):

So just to be sure, x=-1?

OpenStudy (anonymous):

Another way to think of it would be to write out the exponents of each quantity, like the root is to the one half, then if you multiply (1/2)(2), you get (2/2)=1 right? So the exponents cancel but the quantity inside stays the same.

OpenStudy (anonymous):

No, it doesn't equal -1, I missed an addition step.

OpenStudy (anonymous):

The result would be something like -x+3

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

You get -x+3 because if you square everything, you get x-2-x+6-1, so combine like terms.

OpenStudy (anonymous):

It seems right until I plut it into \[\sqrt{2(3)+6}\] That turns into:\[\sqrt{12}\] But there is no complete square root of 12, so whatever it is -1, it can't be right.

OpenStudy (anonymous):

That leaves me with: 1-\[1-\sqrt{12}-1=0\]

OpenStudy (anonymous):

No, that's not it. The result is not a number unless you have something to set it equal to. You get an expression as an answer.

OpenStudy (anonymous):

In the original problem, the -1 was =1 and I moved it to the left side, I just didn't include that step.

OpenStudy (anonymous):

What was the left side originally?

OpenStudy (anonymous):

\[\sqrt{x-2}-\sqrt{2x+6}=1\]

OpenStudy (anonymous):

Either way, if I square the 1 on either side, I still end up with -1.

OpenStudy (anonymous):

No, I mean the full equation. In your question, you gave an expression. An equation must have an equal sign. Was zero on the other side of the equation originally? Or was nothing there?

OpenStudy (anonymous):

The last equation I posted was the full equation.

OpenStudy (anonymous):

Then your answer will be an expression, and you cannot move it to the other side.

OpenStudy (anonymous):

It says I have to solve for x though, I have to create 0 to solve for x.

OpenStudy (anonymous):

Oh so you mean it was given to you already set equal to one? In the problem they gave you?

OpenStudy (anonymous):

Yes, exactly. I just left out that step in my original post. The intent was to save you some work, but I will know better next time :P

OpenStudy (anonymous):

Lol well you have to say that you need to solve for x, otherwise the answer will be an expression. Then here's what you do: \[\sqrt{x-2} - \sqrt{x+6} = 1\] Which is the same as: \[(x-2)^{1/2} - (x+6)^{1/2} = 1\] In order to solve, we must get rid of the root, and we do this by squaring the entire equation. \[(x-2)^{(1/2)(2)} - (x+6)^{(1/2)(2)} = 1^{2}. = x - 2 - x + 6 = 1\]

OpenStudy (anonymous):

Oh! I tried to do that at first, but I didn't think to square those answers.

OpenStudy (anonymous):

Yeah, just do all that, and then solve for x like a normal linear equation.

OpenStudy (anonymous):

Thank you, ahhhh, I was so close :P

OpenStudy (anonymous):

Haha no worries. And I just realized I forgot the 2x so just pretend it's there lol